OpenStudy (anonymous):

Evaluate the following expression for x = 5 and y = -7 5x + 7y – 3x + 8 Using 'Pemdas' i get -47.But thats not a answer in the multiple choice. A. -31 B. -7 C. -16 D. 29

OpenStudy (anonymous):

5x+7y-3x+8 5(5)+7(-7)-3(5)+8=25-49-15+8=-31

OpenStudy (anonymous):

Oh okay, thanks.